Grade 12Integral Calculus

evaluate
∫ sqrt(1-cos2x / 2) dx

taniska
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \sqrt{\frac{1-cos2x}{2}}dx
cos2x = 1 - 2sin^{2}x
\frac{1 - cos2x}{2} = sin^{2}x
I = \int \sqrt{sin^{2}x}dx
I = \int sinx.dx
I = -cosx + constant